Question 22

Stock Y has an expected return of 14% and beta of 1.80. Stock Z has an expected return of 11.50% and beta of 1.10. If the risk-free rate is 3.5% and the market risk premium is 6.5%, which security is overvalued?

Stock Y, because it plots below the SML
Stock Z, because it plots below the SML
Stock Z, because it plots above the SML
Stock Y, because it plots above the SML
No answer text provided.

Flag this Question

Question 23

ABC Inc.'s stock has a 40% chance of producing a 8.6% return, a 25% chance of producing a 11.2% return, and a 35% chance of producing a 21.7% return. What is ABC Inc's expected return?

Enter your answer in percentages rounded off to two decimal points. Do not type % in the answer box.

Flag this Question

Question 24

The common stock of ABC has a beta of 1.22. The market risk premium is 10.7 percent and the risk-free rate is 3.1 percent. What is the firm’s expected return, E(Ri)?

Hint: Use the CAPM equation to get the answer.

Enter your answer in percentages rounded off to two decimal points. Do not enter % in the answer box.

Flag this Question

Question 25

You own a portfolio invested 28.44% in Stock A, 14.81% in Stock B, 27.19% in Stock C, and the remainder in Stock D. The beta of these four stocks are 0.78, 1.14, 0.34, and 1.04. What is the portfolio beta?

Note: Enter your answer rounded off to two decimal points. For example, if your answer is 12.345 then enter as 12.35 in the answer box.
Answer #1
22...According to the CAPM,SML equation is as follows:
Reqd. return on Equity, ke=RFR+Beta*Market risk premium
So, calculating the reqd. return on Stock Y,
3.5%+(1.8*6.5%)=
15.20%
AS the expected return in the market for this stock is 14% , for this measure of beta,it will be plotted below the SML in the graph.
It is overvalued in price because return of 14% does not seem to overcome the risks associated with the stock.
Similarly,, calculating the reqd. return on Stock Z,
3.5%+(1.1*6.5%)=
10.65%
AS the expected return in the market for this stock is 11.5% , for this measure of beta,it will be plotted above the SML in the graph.
Hence Stock Z is under-valued in price because of its greater return of 11.5%   that overcomes the risks associated with the stock.
So, ANSWER is: Security that is overvalued is:
Stock Y, because it plots below the SML
23..ABC's expected reurn =Sum of all probable returns
ie.ABC's expected Return=(40%*8.6%)+(25%*11.2%)+(35%*21.7%)=
13.84%
24…AS per CAPM,
The expected return of ABC=3.1%+(1.22*10.7%)
16.15%
25..The portfolio beta is the weighted beta of its individual components
ie. (28.44%*0.78)+(14.81%*1.14)+(27.19%*0.34)+(29.56%*1.04)=
0.79